A 170 gram ball is dropped from rest off of a cliff. the ball then drops straight down to the ground, 19m below. if the speed of the ball just before it hits the ground is 8 m/s, how much work is done to the ball by the fore of air resistance as it falls?
Answer #1

By work energy theorem Work done by all forces=Change in K.E W +W.= K g f 1-K f 1 =mgh+W, = 3 mo?-5mxo +0.17x9.8x19+W, = 5x0.